# ORION: Option-Regularized Deep Reinforcement Learning for Cooperative Multi-Agent Online Navigation

### v2

@@ Abstract
Existing methods for multi-agent navigation typically assume fully known environments, offering limited support for partially known scenarios such as warehouses or factory floors. There, agents may need to plan trajectories that balance their own path optimality with their ability to collect and share information about the environment that can help their teammates reach their own goals. To these ends, we propose ORION, a novel deep reinforcement learning framework for cooperative multi-agent online navigation in partially known environments. Starting from an imperfect prior map, ORION trains agents to make decentralized decisions, coordinate to reach their individual targets, and actively reduce map uncertainty by sharing online observations in a closed perception–action loop. We first design a shared graph encoder that fuses prior map with online perception into a unified representation, providing robust state embeddings under dynamic map discrepancies. At the core of ORION is an option-critic framework that learns to reason about a set of high-level cooperative modes that translate into sequences of low-level actions, allowing agents to switch between individual navigation and team-level exploration adaptively. We further introduce a dual-stage cooperation strategy that enables agents to assist teammates under map uncertainty, thereby reducing the overall makespan. Across extensive maze-like maps and large-scale warehouse environments, our simulation results show that ORION achieves high-quality, real-time decentralized cooperation over varying team sizes, outperforming state-of-the-art classical and learning-based baselines. Finally, we validate ORION on physical robot teams, demonstrating its robustness and practicality for real-world cooperative navigation.
@@ I Introduction
Multi-agent navigation is a fundamental robotic problem, where a team of agents must reach their individual targets while minimizing the overall makespan, i.e., the travel distance/time until the last agent reaches its assigned target. Beyond the challenges of single-agent navigation, such as long-horizon planning and dynamic obstacle avoidance, multi-agent navigation further introduces team-level cooperation: agents need to anticipate/react to dynamic interactions arising from neighboring robots and changes in shared environments.
Figure 1: Overview of ORION for cooperative multi-agent online navigation. Each agent is assigned a target and navigates over a prior map that may differ from the ground truth. Agents maintain/share (a) prior map, (b) current map, and (c) combined map that fuse prior/online sources to reason about partially changed environments. During navigation, agents not only pursue their own targets but also cooperate by sharing information and assisting others. For example, the red agent reaches its target early and then helps the yellow agent by exploring uncertain regions before returning. ORION enables such adaptive cooperation both before and after arrival on-goal, ultimately reducing the team’s makespan by coordinating agents to contribute where they are most needed in a decentralized way.
@@
However, many real-world deployments, especially in warehouses or factory floors, operate under imperfect/outdated prior maps due to frequent layout changes and occlusions. This motivates a nonstandard yet practical variant of multi-agent navigation, where robots must plan and coordinate online while jointly reducing map uncertainty through decentralized exploration and information sharing. This setting amplifies two key challenges: balancing individual target-reaching with team-level information gathering, and reasoning about dynamic multi-robot interactions under partial observability. In such cases, robots must navigate while continuously updating their belief of the map, requiring richer team-level reasoning that goes beyond classical (sometimes heuristic and offline) path planning. That is, environmental uncertainty forces the team to adaptively trade off between exploiting known areas and exploring unknown areas that may benefit the team’s overall objective (e.g., find shortcuts for other agents, or help them avoid dead-ends). In doing so, cooperation becomes more complex, since a robot may need to temporarily deviate from its shortest path or intentionally gather information that primarily benefits others, sometimes at the expense of its own interests/budget.
@@
Early works in multi-agent navigation primarily focused on ensuring collision-free motion at the kinematic level, typically by decoupling target-driven planning from local obstacle/neighboring robots avoidance [10, 28]. These approaches largely treat interactions among the robot team as safety constraints rather than opportunities to enhance team-level performance: robots rarely leverage team-level coordination and shared global information to guide long-horizon decision-making. Another line of work studies formation-preserving or centroid-driven navigation, where coordinated motions yield more globally consistent behaviors, e.g., maintaining formation structure and avoiding oscillations/conflicts [24, 20]. Yet, these approaches typically assume a known environment or fixed/specific formation objectives, which limits their flexibility in open-ended navigation tasks where robots must dynamically adapt their behaviors to evolving map information. Another closely related domain is Multi-Agent Path Finding (MAPF), where the goal is to compute globally coordinated, conflict-free trajectories for all robots (usually offline), often emphasizing scalability in large structured environments [14, 13, 21]. Despite the global coordination achieved by MAPF approaches, they usually assume a fully known environment and lack the ability to reason about partial observability, exploration, or online map updates.
@@
To address these challenges, we propose ORION, a novel deep reinforcement learning-based framework designed for cooperative multi-agent online navigation. ORION first introduces a fused graph-based representation of each agent’s prior and online updated maps, enabling the downstream policy network to jointly reason about known structures and newly observed areas. We further propose an option-critic structure in which an option denotes a high-level policy that either focuses on the agent’s individual navigation or on supporting teammates by collecting team-level informative observations about the environment. Each option persists for multiple time steps and provides the corresponding low-level action guidance. A learned termination mechanism then decides, based on local observations and the team’s state, whether to continue the current option or switch to a more suitable one. Embedding this option mechanism into both the policy and critic networks enables ORION to judge/evaluate the long-term impact of option sequencing, allowing robots to adopt behaviors that balance individual progress with team-level gains.
Within this option framework, we further design a two-stage cooperation strategy that captures distinct behaviors before and after an agent reaches its target. Before arrival, agents navigate toward their targets while opportunistically gathering information that can reduce future detours and alleviate team-level information gaps. After arrival, instead of remaining idle, agents may choose to explore nearby areas to assist other agents still en route, adaptively deciding how much additional exploration is beneficial without jeopardizing their timeliness. This strategy is validated in our experiments to benefit the makespan by resolving remaining uncertainties for agents still “suffering” from individual exploration-exploitation trade-offs.
We evaluate ORION extensively across hundreds of simulated maps, large-scale Gazebo environments, and real-world deployments with varying team sizes and start–target assignments. ORION consistently outperforms state-of-the-art classical and learning-based planners by $6.9\text{\,}\%-$$13.4\text{\,}\%$ in makespan while maintaining real-time inference efficiency. We further demonstrate up to an additional $14.2\text{\,}\%$ improvement arising specifically from our two-stage cooperation strategy. Finally, real-world experiments further demonstrate that ORION transfers without additional training/tuning, highlighting its practicality for cooperative navigation under partial observability.
@@ II-A Navigation in Partially Known Environments
Navigation in fully known environments has been extensively studied, with search-based planners relying on efficient heuristic-guided graph expansion and sampling-based planners exploring the continuous space through incremental sampling [12, 17, 11]. While these methods are effective in static and fully known settings, they tend to degrade in partially known environments due to their limited ability to anticipate unobserved environment and the heavy computation required for frequent replanning.
@@ II-B Cooperative Multi-Agent Navigation
Beyond single-robot navigation, achieving cooperation across a team of robots poses more challenges. Most works in multi-agent navigation focus on guaranteeing collision-free motion at the kinematic level [10, 28], often in cluttered/crowded environments. These methods primarily treat other agents as local dynamic constraints, and their policies are designed to ensure short-horizon safety and feasibility. While some of them explicitly consider environmental uncertainty [25], this decision-making is still tightly coupled to individual belief updates, making it difficult to endow robots with long-horizon, team-level reasoning in large, partially known environments. Another line of work studies multi-agent navigation while keeping a free-floating formation of specific shape [24, 20]. These approaches design behavior-based rules to induce globally consistent motion, but they typically assume known environments, fixed formation objectives, or pre-specified connectivity patterns.
Multi-Agent Path Finding (MAPF) offers a more global view of coordination, aiming to compute conflict-free trajectories that minimize makespan. Representative methods include bounded-suboptimal [14], prioritized/rule-based search [13], as well as learning-augmented [21] and lifelong variants [8] that scale to thousands of agents. Despite their strong scalability and coordination in discrete, fully known environments, extending MAPF to continuous, partially known environments with online execution remains challenging. A few works move closer to cooperative reasoning by studying multi-agent path topologies without explicit communication [18], enabling targeted rendezvous/message passing among agents [3, 23], or co-optimizing policies with reconfigurable environments [4]. Yet, they remain largely heuristics-based, relying on implicit communication that emerges from local observations rather than deliberate intent sharing, and thus lack mechanisms to model temporally extended, altruistic cooperation, such as when to sacrifice individual progress or re-engage to assist teammates.
Figure 2: Option-regularized policy and multi-agent critic networks. The combined and current encoders fuse prior information with online observations into joint features. A termination head and option decoder then decide whether to maintain the current option or switch to a new valid one, while the waypoint decoder integrates the option feature with the current node feature to select a waypoint from the agent’s neighboring nodes. In parallel, the critic network, conditioned on the states, actions, and selected options of all agents, provides centralized value estimates. During training, it leverages the ground-truth map as privileged information to estimate option–state values that capture the long-term team return.
@@ III-A Problem Formulation
We study multi-agent navigation in partially known environments with online map sharing. A team of $N$ agents $\mathcal{A}=\{1,\dots,N\}$ is initially provided a 2D prior occupancy map $\mathcal{M}^{-}$ consisting of free regions $\mathcal{M}_{f}^{-}$ and occupied regions $\mathcal{M}_{o}^{-}$. The true environment is described by a ground truth map $\mathcal{M}_{g}$, which may partially differ from $\mathcal{M}^{-}$ due to layout changes, such as shelf rearrangements. During execution, agents share sensor observations through perfect global communication to incrementally build a current map $\mathcal{M}$, which is initially unknown and gradually reveals the true structure of the environment. To reason jointly over prior knowledge and newly revealed areas, agents maintain a combined map $\hat{\mathcal{M}}$, initialized as $\mathcal{M}^{-}$ and continuously updated using observations accumulated in $\mathcal{M}$.
The environment is modeled as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where nodes $v\in\mathcal{V}$ are uniformly sampled from free space in $\hat{\mathcal{M}}$ (and $\mathcal{M}$) and edges $(v_{i},v_{j})\in\mathcal{E}$ represent collision-free transitions between neighboring nodes, with each node connected to up to $k$ neighbors. Each agent starts from an initial node $s_{i}\in\mathcal{V}$ and aims to reach a designated target node $g_{i}\in\mathcal{V}$. At each decision step, an agent may move to a neighboring node or remain stationary. Agents must avoid vertex and edge collisions, i.e., no two agents may occupy the same node at the same time, i.e., $v^{i}_{t}\neq v^{i}_{t}$ for all $i\neq j$ or traverse the same edge in opposite directions, i.e., $(v^{i}_{t},v^{i}_{t+1})\neq(v^{j}_{t+1},v^{j}_{t})$ for all $i\neq j$. The objective is to generate a set of collision-free trajectories $\Psi=\{\psi_{1},\dots,\psi_{N}\}$ that guide all agents to their targets through sequence of neighboring nodes while minimizing the team makespan, defined as $\min_{\Psi}\max_{i\in\mathcal{A}}T_{i}$, where $T_{i}$ denotes the arrival time (or distance) of agent $i$.
@@ III-B Observation as Informative Graph
Each node/viewpoint is augmented with semantic attributes beyond 2D coordinates to obtain richer representations: (1) Prior utility $u_{p,j}$: extending the notion of frontiers [27], we define prior frontiers $\mathcal{F}_{p}$ as boundaries between free space and uncertain obstacles, and let $u_{p,j}$ denote the number of such prior frontiers that are visible, i.e., along a collision-free line of sight, from $\hat{v}_{j}\in\hat{V}_{t}$ within a sensing range $r_{\rm fov}$. (2) Utility $u_{j}$: defines the same quantity but computes it on the current map $\mathcal{M}$. (3) Visited flag $\delta_{i,j}$: indicates whether agent $i$ has visited $v_{j}$. (4) Verified signal $s_{j}$: marks whether $v_{j}$ lies within $\mathcal{M}$. (5) Occupancy indicator $p_{j}$: encodes, as a ternary variable, whether $v_{j}$ is currently occupied by agent $i$, by another agent, or by no agent. (6) Target binary $t_{j}$: specifies whether $v_{j}$ coincides with agent $i$’s own target or with another agent’s target (if any). We further cluster all viewpoints with non-zero utility within a range $r_{b}$ into a beacon set [15], representing candidate regions to navigate. We further construct a planning graph $\bar{G}_{t}=(\bar{V}_{t},\bar{E}_{t})$ on $\mathcal{M}$, containing samples from both free and unknown cells, where edges connect collision-free pairs of neighbors. These structures allow us to define: (7) Navigation guidepost $g_{n,j}$: the Dijkstra trajectory on $\bar{G}_{t}$ from $v_{j}$ to the beacon nearest the agent’s assigned target. (8) Cooperation guidepost $g_{c,j}$: if any target remains unverified, i.e., not in the explored area, we compute the trajectory from $v_{j}$ to the beacon nearest the closest unverified target; otherwise $g_{c,j}=g_{n,j}$ if no targets remain.
III-C Dual-Stage Cooperation Strategy
III-D Policy Networks
III-D1 Graph-based Encoders

### v3

@@ Abstract
Existing methods for multi-agent navigation typically assume fully known environments, offering limited support for partially known scenarios with outdated or imperfect prior maps, such as warehouses or factory floors. There, agents need to balance path optimality with collecting and sharing environmental information to help teammates reach their own targets. To these ends, we propose ORION, a novel deep reinforcement learning framework for cooperative multi-agent online navigation in partially known environments. Starting from an imperfect prior map, ORION trains agents to make decentralized decisions, coordinate toward individual targets, and actively reduce task-relevant map uncertainty through online observation sharing in a closed perception–-action loop. We first design a shared graph encoder that fuses prior map with online perception into a unified representation, providing robust state embeddings under environmental discrepancies. At the core of ORION is an option-critic framework that learns high-level cooperative modes translated into sequences of low-level actions, enabling adaptive switching between individual navigation and team-level exploration. We further introduce a dual-stage cooperation strategy that allows agents to assist teammates under map uncertainty, thereby reducing the overall makespan. Across extensive maze-like maps and large-scale warehouse environments, ORION achieves high-quality real-time decentralized cooperation while scaling to up to 10 robots, outperforming state-of-the-art classical and learning-based baselines. Finally, we validate ORION on physical robot teams, demonstrating its robustness and practicality for real-world cooperative navigation.
@@ I Introduction
Multi-agent navigation is a fundamental robotic problem, where a team of agents must reach individual targets while minimizing the overall makespan, i.e., the travel distance or time until the last agent reaches its assigned target. Beyond the challenges of single-agent navigation, such as long-horizon planning and dynamic obstacle avoidance, multi-agent navigation further introduces team-level cooperation: agents need to anticipate/react to dynamic interactions arising from neighboring robots and changes in shared environments. However, many real-world deployments, especially in warehouses or factory floors, operate under imperfect or outdated prior maps due to frequent layout changes and occlusions. This motivates a nonstandard yet practical variant of multi-agent navigation, where robots must plan and coordinate online while jointly reducing map uncertainty through decentralized exploration and information sharing. This setting amplifies two key challenges: balancing individual target-reaching with team-level information gathering, and reasoning about dynamic multi-robot interactions under partial observability. In such cases, robots must navigate while continuously updating their belief of the map, requiring richer team-level reasoning that goes beyond classical (sometimes heuristic and offline) path planning. That is, environmental uncertainty forces the team to adaptively trade off between exploiting known areas and exploring unknown areas that may benefit the team’s overall objective (e.g., find shortcuts for other agents, or help them avoid dead-ends). Cooperation is therefore more complex: a robot may need to temporarily deviate from its shortest path or intentionally gather information that primarily benefits others, sometimes at the expense of its own budget.
Figure 1: Overview of Multi-Agent Online Navigation. Each agent is assigned a target and navigates over a prior map that may differ from the ground truth. Agents maintain/share (a) prior map, (b) current map, and (c) combined map that fuse prior/online sources to reason about partially changed environments. During navigation, agents not only pursue their own targets but also cooperate by sharing information and assisting others. For example, the red agent reaches its target early and then helps the yellow agent by exploring uncertain regions before returning. ORION enables such adaptive cooperation both before and after arrival on-goal, ultimately reducing the team’s makespan by coordinating agents to contribute where they are most needed in a decentralized way.
@@
Early works in multi-agent navigation primarily focused on ensuring collision-free motion at the kinematic level, typically by decoupling target-driven planning from local obstacle/neighboring robots avoidance [13, 32, 29]. These approaches largely treat interactions among the robot team as safety constraints rather than opportunities to enhance team-level performance: robots rarely leverage team-level coordination and shared global information to guide long-horizon decision-making. Another line of work studies formation-preserving or centroid-driven navigation, where coordinated motions yield more globally consistent behaviors, e.g., maintaining formation structure and avoiding oscillations/conflicts [27, 23]. Yet, these approaches typically assume a known environment or fixed/specific formation objectives, which limits their flexibility in open-ended navigation tasks where robots must dynamically adapt their behaviors to evolving map information. Another closely related domain is Multi-Agent Path Finding (MAPF), where the goal is to compute globally coordinated, conflict-free trajectories for all robots (usually offline), often emphasizing scalability in large structured environments [17, 16, 24]. Despite the global coordination achieved by MAPF approaches, they usually assume a fully known environment and lack the ability to reason about partial observability, exploration, or online map updates.
@@
To address these challenges, we propose ORION, a novel deep reinforcement learning-based framework designed for cooperative multi-agent online navigation. ORION first introduces a fused graph-based representation of each agent’s prior and online updated maps, enabling the downstream policy network to jointly reason about known structures and newly observed areas. We further propose an option-critic structure in which an option denotes a high-level policy that either focuses on the agent’s individual navigation or on supporting teammates by collecting team-level informative observations about the environment. Each option persists for multiple time steps and provides the corresponding low-level action guidance. A learned termination mechanism then decides, based on local observations and the team’s state, whether to continue the current option or switch to a more suitable one. Embedding this option mechanism into both the policy and critic networks enables ORION to judge/evaluate the long-term impact of option sequencing, allowing robots to adopt behaviors that balance individual progress with team-level gains. Within this option framework, we further design a two-stage cooperation strategy that captures distinct behaviors before and after an agent reaches its target. Before arrival, agents navigate toward their targets while opportunistically gathering information that can reduce future detours and alleviate team-level information gaps. After arrival, instead of remaining idle, agents may choose to explore nearby areas to assist other agents still en route, adaptively deciding how much additional exploration is beneficial without jeopardizing their timeliness. This strategy is validated in our experiments to benefit the makespan by resolving remaining uncertainties for agents still “suffering” from individual exploration-exploitation trade-offs. We evaluate ORION extensively across hundreds of simulated maps, large-scale Gazebo environments, and real-world deployments with varying team sizes and start–target assignments. ORION consistently outperforms state-of-the-art classical and learning-based planners by $6.9\text{\,}\%-$$13.4\text{\,}\%$ in makespan while maintaining real-time inference efficiency. We further demonstrate up to an additional $14.2\text{\,}\%$ improvement arising specifically from our two-stage cooperation strategy. Finally, real-world experiments further demonstrate that ORION transfers without additional training/tuning, highlighting its practicality for cooperative navigation under partial observability.
@@ II-A Navigation in Partially Known Environments
Navigation in fully known environments has been extensively studied, with search-based planners relying on efficient heuristic-guided graph expansion and sampling-based planners exploring the continuous space through incremental sampling [15, 20, 14, 6]. While these methods are effective in static and fully known settings, they tend to degrade in partially known environments due to their limited ability to anticipate unobserved environment and the high computational cost of frequent replanning.
@@ II-B Cooperative Multi-Agent Navigation
Beyond single-robot navigation, achieving cooperation across a team of robots poses more challenges. Most works in multi-agent navigation focus on guaranteeing collision-free motion at the kinematic level [13, 32, 29], often in cluttered/crowded environments. These methods primarily treat other agents as local dynamic constraints, and their policies are designed to ensure short-horizon safety and feasibility. While some of them explicitly consider environmental uncertainty [28], this decision-making is still tightly coupled to individual belief updates, making it difficult to endow robots with long-horizon, team-level reasoning in large, partially known environments. Another line of work studies multi-agent navigation while keeping a free-floating formation of specific shape [27, 23]. These approaches design behavior-based rules to induce globally consistent motion, but they typically assume known environments, fixed formation objectives, or pre-specified connectivity patterns. Multi-Agent Path Finding (MAPF) offers a more global view of coordination, aiming to compute conflict-free trajectories that minimize makespan. Representative methods include bounded-suboptimal [17], prioritized/rule-based search [16], as well as learning-augmented [24] and lifelong variants [11] that scale to thousands of agents. Despite their strong scalability and coordination in discrete, fully known environments, extending MAPF to continuous, partially known environments with online execution remains challenging. A few works move closer to cooperative reasoning by studying multi-agent path topologies without explicit communication [21], enabling targeted rendezvous/message passing among agents [4, 26, 1], or co-optimizing policies with reconfigurable environments [7]. Yet, they remain largely heuristics-based, relying on implicit communication that emerges from local observations rather than deliberate intent sharing, and thus lack mechanisms to model temporally extended, altruistic cooperation, such as when to sacrifice individual progress or re-engage to assist teammates.
Figure 2: Policy and multi-agent critic networks of ORION. The combined and current encoders fuse prior information with online observations into joint features. A termination head and option decoder then decide whether to maintain the current option or switch to a new valid one, while the waypoint decoder integrates the option feature with the current node feature to select a waypoint from the agent’s neighboring nodes. In parallel, the critic network, conditioned on the states, actions, and selected options of all agents, provides centralized value estimates. During training, it leverages the ground-truth map as privileged information to estimate option–state values that capture the long-term team return.
@@ III-A Problem Formulation
We study multi-agent navigation in partially known environments with online map sharing. A team of $N$ agents $\mathcal{A}=\{1,\dots,N\}$ is initially provided a 2D prior occupancy map $\mathcal{M}^{-}$ consisting of free regions $\mathcal{M}_{f}^{-}$ and occupied regions $\mathcal{M}_{o}^{-}$. The true environment is described by a ground truth map $\mathcal{M}_{g}$, which may partially differ from $\mathcal{M}^{-}$ due to layout changes, such as shelf rearrangements. During execution, agents share sensor observations through perfect global communication to incrementally build a current map $\mathcal{M}$, which is initially unknown and gradually reveals the true structure of the environment. To reason jointly over prior knowledge and newly revealed areas, agents maintain a combined map $\hat{\mathcal{M}}$, initialized as $\mathcal{M}^{-}$ and continuously updated using observations accumulated in $\mathcal{M}$. The environment is modeled as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where nodes $v\in\mathcal{V}$ are uniformly sampled from free space in $\hat{\mathcal{M}}$ (and $\mathcal{M}$) and edges $(v_{i},v_{j})\in\mathcal{E}$ represent collision-free transitions between neighboring nodes, with each node connected to up to $k$ neighbors. Each agent starts from an initial node $s_{i}\in\mathcal{V}$ and aims to reach a designated target node $g_{i}\in\mathcal{V}$. At each decision step, an agent may move to a neighboring node or remain stationary. Agents must avoid vertex and edge collisions, i.e., no two agents may occupy the same node at the same time, i.e., $v^{i}_{t}\neq v^{i}_{t}$ for all $i\neq j$ or traverse the same edge in opposite directions, i.e., $(v^{i}_{t},v^{i}_{t+1})\neq(v^{j}_{t+1},v^{j}_{t})$ for all $i\neq j$. The objective is to generate a set of collision-free trajectories $\Psi=\{\psi_{1},\dots,\psi_{N}\}$ that guide all agents to their targets through sequence of neighboring nodes while minimizing the team makespan, defined as $\min_{\Psi}\max_{i\in\mathcal{A}}T_{i}$, where $T_{i}$ denotes the arrival time (or distance) of agent $i$.
@@ III-B Observation as Informative Graph
Each node/viewpoint is augmented with semantic attributes beyond 2D coordinates to obtain richer representations: (1) Utility $u_{j}$: Following [31, 18], frontier measures as boundaries between free space and unknown areas in the current map $\mathcal{M}$, and the utility $u_{j}$ denotes the number of such frontiers at node $v_{j}$, which are visible, i.e., along a collision-free line of sight. (2) Prior utility $u^{p}_{j}$: Similar to frontiers, prior frontiers represent the boundary between confirmed free space (validated by the current map) and uncertain obstacles inferred from the prior map. It combines confirmed observations in $\mathcal{M}$ with prior knowledge in $\mathcal{M}^{-}$. These prior frontiers capture the potential information gain suggested by the prior map, offering a long-range heuristic of current unknown areas. Prior utility $u^{p}_{j}$ is defined as the number of observable prior frontiers at node $v_{j}$. (3) Visited flag $\delta_{j,i}$: indicates whether $v_{j}$ has been visited by agent $i$. (4) Verified signal $s_{j}$: marks whether $v_{j}$ lies within $\mathcal{M}$. (5) Occupancy indicator $p_{j,i}$: as a ternary variable, encodes whether $v_{j}$ is currently occupied by agent $i$. (6) Target binary $t_{j,i}$: specifies whether $v_{j}$ coincides with agent $i$’s own target or with another agent’s target (if any). We further cluster all viewpoints with non-zero utility within a range $r_{b}$ into a beacon set [18], representing candidate regions to navigate. We further construct a planning graph $\bar{G}_{t}=(\bar{V}_{t},\bar{E}_{t})$ on $\mathcal{M}$, containing samples from both free and unknown cells, where edges connect collision-free pairs of neighbors. Based on $\mathcal{M}$, we define two guideposts that facilitate coordinated navigation: (7) Navigation guidepost $g_{i}$: the Dijkstra trajectory on $\bar{G}_{t}$ from agent $i$ to the beacon nearest the agent’s assigned target [5], which serves as a feasible trajectory cue for self-navigation. (8) Cooperation guidepost $g^{c}_{i}$: if any other target remains unverified, i.e., not in the explored area, we compute the trajectory from agent $i$ to the beacon nearest the closest unverified target; otherwise $g_{i}=g^{c}_{i}$ if all targets are verified. This guidance makes it feasible for agents to assist teammates.
III-C Policy and Critic Networks
III-C1 Graph-based Encoders